OpenStudy (anonymous):

Medal <3 which ratios form a proportion? A. 3/15 and 12/55 B. 8/24 and 12/35 C. 5/18 and 25/90 D. 4/11 and 16/25

OpenStudy (yanasidlinskiy):

Cross multiply and see if they equal the same on each side.

OpenStudy (anonymous):

alternatively reduce each fraction to lowest terms and see which ones are equal

OpenStudy (anonymous):

cheating alternative use a calculator see which ones give the same decimal

OpenStudy (anonymous):

I don't get it, so i have to divide

OpenStudy (anonymous):

1/ Multiply top and bottom by one correct number and see if you get the second number: 3/15 : 12/55 No. 3x4= 12 but 15x4 does not equal 55 ....................... 2/ 9/15 x2 = 18/30... no 36/50 / 4 = 9/15... yes 4/ 2/8= x/48 crossmultiply 8x = 96 x= 96/8 = 12 3/ 2.2x2.5 = 4/ 2/8 = x/48 cross multiply 8x = 96 x = 12 ................... 5/ the same as 4 6/ like #1 7/ 10/750 = 14/x = Do like #4

OpenStudy (anonymous):

a.. the answer is a

OpenStudy (anonymous):

real cheating alternative pick C, it is always C

OpenStudy (anonymous):

what? its not always c

OpenStudy (anonymous):

Is it A or C

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

i used my textbook. that person used a "cheat"

OpenStudy (anonymous):

\[\frac{25}{90}=\frac{5\times 5}{5\times 18}=\frac{5}{18}\]

OpenStudy (anonymous):

So its A

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Don't give out answers, you're helping them cheat without learning anything.

OpenStudy (anonymous):

can i have a medal?

OpenStudy (anonymous):

But what if it not A. I want to get more comments.

OpenStudy (anonymous):

sure have a medal but it is definitely not A

OpenStudy (uri):

Always try solving C. Answer is always C. xD

OpenStudy (anonymous):

So its C

OpenStudy (anonymous):

No giving out answers

OpenStudy (anonymous):

ok. ive had this question before so i know its a. anyway if it ends up being a can i have the medal?

OpenStudy (anonymous):

Yes its C

OpenStudy (anonymous):

\[\frac{12}{55}=\frac{3\times 4}{5\times 11}\] it does not reduce

OpenStudy (anonymous):

ITS C.

OpenStudy (anonymous):

I took the test already, its C.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

if it ends up being A, i will eat my bicycle in macy's window on black friday

OpenStudy (anonymous):

Ok Thanks!

OpenStudy (anonymous):

what happened to no giving out answers? xD

OpenStudy (anonymous):

lol

OpenStudy (uri):

@satellite73 get me something on black friday.

OpenStudy (anonymous):

M too

OpenStudy (anonymous):

Me*